graph the equation using the slope and the y intercept of y=1/5x+7. i know the first corrdinates will be (0,7) need help with finiding the other?
Thanks

Answer by stanbon(57377)

graph the equation using the slope and the y intercept of y=(1/5)x+7. i know the first cordinates will be (0,7) need help with finding the other?
-----
Plot that y-intercept: (0,7)
----
Slope = 1/5 which means "y increases 1/5 when x increases 1".
---
Let x increase by 5, then y will increase by 1+7 = 8.
That gives you the point (5,8)
---
Plot it and draw a line thru the two points you have plotted.
---------------------
Cheers,
Stan H.